Use bond energies as found in Table 9.1 of Tro, Fridgen and Shaw and estimate the energy change ∆ H o (in kJ/mol) for the following reaction:

H—C ≡ C—H + 2 I 2 → CHI 2 —C HI 2

The C-I bond energy is 213 kJ/mol.
Answer #1

Hi,

Hope you are doing well.


__________________________________________________________________________

SOLUTION:

For the given reaction:

H-C=C - H+2(1 - 1) + CHI2 - CHI,

Bond

Type:

No.of

Bonds

Bond

energy:

(kJ/mol)

C - H 4 411
C= C 1 835
I-I 2 148
C - C 1 346
C - 1 4 213

We know that, \Delta H^{o} of this reaction is given by,

ΔΗ° = ΣΒ.Ε(reactants) – ΣΒ.Ε(products)

AH° = [2.B.E(C-H) + B.E(C = H) +2.B.E(I - 1)]–[B.E(C-C)+ 4.B.E(C-1) + 2B.E(C – H)]

.

AH = [2.(411) + (835) +2.(148)] kJ/mol-[(346)+4.(213)+2(411)] kJ/mol

.

:. AH = 1953 kJ/mol - 2020 kJ/mol

.

:: AH = -67 kJ/mol


__________________________________________________________________________

Hope this helped for your studies. Keep learning. Have a good day.

Feel free to clear any doubts at the comment section.


Please don't forget to give a thumps up.

Thank you. :)